OpenStudy (anonymous):

if 4x^2y^2+8x^3y^5 over 2xy^2 is completely simplified to 2xayb + 4xcyd, where a, b, c, and d represent integer exponents, what is the value of c?

OpenStudy (ranga):

Shouldn't it be 2x^ay^b + 4x^cy^d ?

OpenStudy (anonymous):

\(\Large \frac{4x^2y^2+8x^3y^5}{2xy^2}=\frac{4x^2y^2}{2xy^2}+\frac{8x^3y^5}{2xy^2}=2x+4x^2y^3\) Thus, \(\Large a=1, b=0,c=2,d=3 \)

OpenStudy (anonymous):

\(\Large y^0=1\)

OpenStudy (anonymous):

@ranga no the equation itself is correct, I copied and pasted it.

OpenStudy (ranga):

The equation is correct but I think the part "simplified to 2xayb + 4xcyd" should be: "simplified to 2x^ay^b + 4x^cy^d". (the ^s are missing).

OpenStudy (anonymous):

ooh I see what your saying, yes your right :3